# Why Mask Diffusion does not Work

### v1

@@ 1 Introduction
To adapt diffusion for discrete language data, one of the primary directions has been the development of continuous-space models (li2022diffusionlm, ; dieleman2022continuous, ; han2022ssdlm, ). This approach first maps discrete tokens into a continuous embedding space where the denoising process is performed. The forward process gradually transforms the token embeddings into noise, and the reverse process learns to invert this corruption, starting from noise and progressively generating a clean embedding. After a denoised embedding is generated, a final rounding step maps it back to a discrete token from the vocabulary.
In recent years, discrete-space DLMs have gained increasing popularity as they define the diffusion process directly on the vocabulary of tokens, avoiding the need for a continuous embedding space during diffusion itself. In this paradigm, the forward process corrupts a sequence by applying a transition matrix at each step, which defines the probability of a token transitioning to any other token (austin2021structured, ; lou2024discrete, ; hoogeboom2021argmax, ). The model then learns to reverse these transitions by predicting the original token distribution given the corrupted sequence.
Masked Diffusion Models represent a modern and highly effective evolution of discrete diffusion language models, forming the foundation for several recent large-scale efforts such as DiffuLLaMA, LLaDA and Dream (gong2025scaling, ; nie2025large, ; ye2025dream, ). The core principle involves a forward process where tokens in a sequence are probabilistically replaced with a special [MASK] token (austin2021structured, ). During training, the model learns to predict the original content for these [MASK] positions, with the loss computed only over the [MASK] tokens. The generation process starts with a fully [MASK] sequence. In each iterative step, the model predicts several tokens. This iterative refinement continues until all [MASK] tokens are resolved.
@@ 2.2 Mask Diffusion Modeling
where $\langle f_{\theta}(x_{t}),m\rangle=0$ (shi2024simplified, ).
@@
In the continuous time limit where $T\to\infty$ (shi2024simplified, ), we set a small timestep $\Delta_{t}=t(i)-s(i)=\frac{1}{T}\in(0,1)$. The sum over timesteps becomes an integral, and we have a limit
@@ 3.2.1 Weighted Loss Function
The loss function (Equation 13) essentially measures the model’s ability to reconstruct the original data from the noised input at any time step $t$. The only explicit dependence on $t$ appears in the weighting term (e.g., $\frac{\alpha_{t^{\prime}}}{1-\alpha_{t}}$). Theoretically, however, if the model is sufficiently trained (all loss equal to 0), these coefficients are not crucial. The role of the weights is mainly to encourage the model to focus more on cases with small $t$, while down-weighting the loss at large $t$, thereby stabilizing the training curve(shi2024simplified, ).
@@ 3.2.2 Model Output as Marginal Distributions
In discrete settings, it is necessary to manually define neighborhoods, as in concrete score matching(meng2022concrete, ). Mask diffusion is essentially a single-token transition process, similar to SEDD(lou2024discrete, ). Since noise is added independently to each token, The forward process can be regarded as many independent single-token transitions executed in parallel. This, in turn, implies that the reverse denoising process is also fundamentally a single-token transition problem: the model learns how to update one token at a time. However, when multiple tokens are updated simultaneously, there is no guarantee of their mutual coherence. We will demonstrate it in the following example.
@@
In the above experiment, the model fails to retain the joint probability of the first token pair. Instead, it only ouput $P(A|CD)=0.55$ and $P(A^{\prime}|CD)=0.45$ at the first [MASK] position and $P(B|CD)=0.69$ and $P(B^{\prime}|CD)=0.31$ at the second [MASK] position (Under Assumption 1, we omit the explicit form of the [MASK] tokens when computing conditional probabilities). If now you decide to predict the first two tokens at once, your choice will be $AB$ while $A^{\prime}B$ is the most probable pair.
@@ 3.2.4 Experiment
Our estimation follows the same trend as the actual behavior in the beginnig, however, the true average maximum probability usually converges to a smaller value (but still noticeably greater than zero). Our predicted curve does not rise afterwards, since our assumption does not account for length constraints or the special end-of-text token.
@@ 3.3 Parallel Sampling
The $PPL$ of a senquence $(x^{1},x^{2},\cdots,x^{n})$ is defined as $P(x^{1},x^{2},\cdots,x^{n})^{-\frac{1}{n}}$.
@@
Although mask diffusion also learns single-token transitions, it differs from SEDD in that the denoising process fixes the unmasked tokens. While certain remasking techniques exist(wang2025remasking, ), they can not achieve parallelism, as remasking multiple unmasked tokens simultaneously may cause substantial information loss, nor acceleration, since they require an additional prediction step.
@@ 3.4.2 Training Strategies
During inference, we define a small block (e.g., of size 4 or 8) and generate within it using a diffusion process (or parallel sampling with the block size as the upper bound). Once the predictions for the block are completed, we proceed to the next small block, until the model outputs the end-of-text token. This essentially corresponds to semi-autoregression(han2022ssdlm, ), but with smaller blocks; the reason is that [MASK] tokens far from the given condition are difficult to exploit, making large block sizes nearly ineffective.

### v2

@@ 1 Introduction
To adapt diffusion for discrete language data, one of the primary directions has been the development of continuous-space models [li2022diffusionlm, dieleman2022continuous, han2023ssdlm]. This approach first maps discrete tokens into a continuous embedding space where the denoising process is performed. The forward process gradually transforms the token embeddings into noise, and the reverse process learns to invert this corruption, starting from noise and progressively generating a clean embedding. After a denoised embedding is generated, a final rounding step maps it back to a discrete token from the vocabulary.
In recent years, discrete-space DLMs have gained increasing popularity as they define the diffusion process directly on the vocabulary of tokens, avoiding the need for a continuous embedding space during diffusion itself. In this paradigm, the forward process corrupts a sequence by applying a transition matrix at each step, which defines the probability of a token transitioning to any other token [austin2023structured, lou2024discrete, hoogeboom2021argmax]. The model then learns to reverse these transitions by predicting the original token distribution given the corrupted sequence.
Masked Diffusion Models represent a modern and highly effective evolution of discrete diffusion language models, forming the foundation for several recent large-scale efforts such as DiffuLLaMA, LLaDA and Dream [gong2025scaling, nie2025large, ye2025dream]. The core principle involves a forward process where tokens in a sequence are probabilistically replaced with a special [MASK] token [austin2023structured]. During training, the model learns to predict the original content for these [MASK] positions, with the loss computed only over the [MASK] tokens. The generation process starts with a fully [MASK] sequence. In each iterative step, the model predicts several tokens. This iterative refinement continues until all [MASK] tokens are resolved.
@@ 2.2 Mask Diffusion Modeling
where $\langle f_{\theta}(x_{t}),m\rangle=0$ [shi2025simplified].
@@
In the continuous time limit where $T\to\infty$ [shi2025simplified], we set a small timestep $\Delta_{t}=t(i)-s(i)=\frac{1}{T}\in(0,1)$. The sum over timesteps becomes an integral, and we have a limit
@@ 3.2.1 Weighted Loss Function
The loss function (Equation 13) essentially measures the model’s ability to reconstruct the original data from the noised input at any time step $t$. The only explicit dependence on $t$ appears in the weighting term (e.g., $\frac{\alpha_{t^{\prime}}}{1-\alpha_{t}}$). Theoretically, however, if the model is sufficiently trained (all loss equal to 0), these coefficients are not crucial. The role of the weights is mainly to encourage the model to focus more on cases with small $t$, while down-weighting the loss at large $t$, thereby stabilizing the training curve[shi2025simplified].
@@ 3.2.2 Model Output as Marginal Distributions
In discrete settings, it is necessary to manually define neighborhoods, as in concrete score matching[meng2023concrete]. Mask diffusion is essentially a single-token transition process, similar to SEDD[lou2024discrete]. Since noise is added independently to each token, The forward process can be regarded as many independent single-token transitions executed in parallel. This, in turn, implies that the reverse denoising process is also fundamentally a single-token transition problem: the model learns how to update one token at a time. However, when multiple tokens are updated simultaneously, there is no guarantee of their mutual coherence. We will demonstrate it in the following example.
@@
In the above experiment, the model fails to retain the joint probability of the first token pair. Instead, it only output $P(A|CD)=0.55$ and $P(A^{\prime}|CD)=0.45$ at the first [MASK] position and $P(B|CD)=0.69$ and $P(B^{\prime}|CD)=0.31$ at the second [MASK] position (Under Assumption 1, we omit the explicit form of the [MASK] tokens when computing conditional probabilities). If now you decide to predict the first two tokens at once, your choice will be $AB$ while $A^{\prime}B$ is the most probable pair.
@@ 3.2.4 Experiment
Our estimation follows the same trend as the actual behavior in the beginning, however, the true average maximum probability usually converges to a smaller value (but still noticeably greater than zero). Our predicted curve does not rise afterwards, since our assumption does not account for length constraints or the special end-of-text token.
@@ 3.3 Parallel Sampling
The $PPL$ of a sequence $(x^{1},x^{2},\cdots,x^{n})$ is defined as $P(x^{1},x^{2},\cdots,x^{n})^{-\frac{1}{n}}$.
@@
Although mask diffusion also learns single-token transitions, it differs from SEDD in that the denoising process fixes the unmasked tokens. While certain remasking techniques exist[wang2025remasking], they can not achieve parallelism, as remasking multiple unmasked tokens simultaneously may cause substantial information loss, nor acceleration, since they require an additional prediction step.
@@ 3.4.2 Training Strategies
During inference, we define a small block (e.g., of size 4 or 8) and generate within it using a diffusion process (or parallel sampling with the block size as the upper bound). Once the predictions for the block are completed, we proceed to the next small block, until the model outputs the end-of-text token. This essentially corresponds to semi-autoregression[han2023ssdlm], but with smaller blocks; the reason is that [MASK] tokens far from the given condition are difficult to exploit, making large block sizes nearly ineffective.